On July 1, 2020, Coronado Industries purchased machinery for $248000. Salvage value was estimated to be $9500. The machinery will be depreciated over ten years using the double-declining balance method. If depreciation is computed on the basis of the nearest full month, Coronado should record depreciation expense for 2021 on this machinery of of

$39680.

$44640.

$33480.

$45115.
Answer #1

Double declining rate = 200/Useful life = 200/10 = 20%

Depreciation for first year

= 248,000*20%*6/12 = 24,800

Depreciation for second year

= (248,000-24,800)*20%

= 44,640